OpenStudy (anonymous):

Evaluate each polynomial for x = 3 and for x = −3. and dont forget to do the same for x = –3 −x2 − 3x3 − x4

OpenStudy (diyadiya):

x-3 -9-81-81 = -153 x=-3 -9+81-81 = -9

OpenStudy (anonymous):

ty thanks

OpenStudy (anonymous):

thts the one i needed got x=3